# Non-unitary enhanced transfer efficiency in quantum walk search on complex networks

### v2

@@ III.1.2 Numerical results
Tadpole $T_{M,N}$ $M$+$N\!=\!32$+$32$ 0.0317 cycle 0.0317 $48\!=\!N+\lfloor M/2\rfloor=\mathcal{O}(N$+$M)$
@@
Maze $M_{N}$ $N=73$ 0.0138 exit 0.0273 34
Figure 3: Instances $L_{8,8}$ and $T_{8,8}$ of the lollipop (upper) and tadpole (lower) graphs, that are respectively a fusion between the complete graph $K_{8}$ or the cycle graph $C_{8}$ with the path graph $P_{8}$. We respectively refer to the locations of the red, green and blue vertices as complete, shared and path vertices for the lollipop graph, and cycle, shared and path for the tadpole graph.
@@ III.2 Relation to entropy
We now relate the optimal interpolation regime $\omega$ to the evolution of the Von Neumann entropy of the system:
@@
We observe for every run of the SQWS that the entropy first increases up to a maximum, and then decreases until it converges to zero. We also observe that when an increase in the maximum value of $\gamma$ increases the performance of the SQWS, i.e. the transfer to the sink, this translates into a reduction in the time needed to reach the entropy maximum and a faster convergence to zero thereafter. Furthermore, we observe that the optimal interpolation regime $\omega$ is that whose entropy converges to zero the fastest. The presence of the sink vertex $\phi$ introduces dissipation in the system, therefore, as $t\rightarrow\infty$ the walker will end up in the sink with a probability of 1. Thus, the state of the walker converges from $\rho(0)$ to the projector |$\ket{\phi}\bra{\phi}$ |. Although zero entropy indicates a pure state, it does not guarantee that this state is the state towards which the system converges. Therefore, we also compute the $l_{1}$-norm coherence, which is the sum of the off-diagonal elements of $\rho$:
@@
As the initial state is the uniform superposition over the set of vertices $V$, the initial value is $C_{l_{1}}(\rho(0))=|V|-1$. For all graphs on which we run the SQWS, we observe that this quantity decreases monotonically and it converges to zero. Thus, the convergence of entropy to zero does indicate that the walker state is getting closer and closer to the state |$\ket{\phi}\bra{\phi}$ |, as $\rho$ becomes increasingly diagonal over time. We illustrate the entropy evolution over time in Fig. 4 for the complete, the hypercube, the cycle and the lollipop (for the path vertex) graphs.
Looking at Fig. 4, we can see that the choice of $\gamma$ has an impact on the evolution of entropy. For the complete graph on Fig. 4(a), we can see that for $\gamma=0.05$ the entropy that decreases most rapidly is obtained for $\omega=1$, which can be verified with Fig. 2. Furthermore, we observe that for $\omega=0$ entropy increases only very slightly and does not decrease. This is simply because the chosen evolution time is not long enough to observe this phenomenon. We also observe that the $\omega=0.1$ regime is the worst for all choices of $\gamma$ except for $\gamma=0.05$ where the quantum case is the worst. We can see that entropy decreases most rapidly in the quantum case, i.e. $\omega=0$, when $\gamma=2$. This can also be seen in Fig. 2, as we can see that for the $\gamma$ values selected, the best transfer efficiency for the complete graph is obtained for $\gamma=2$. For the hypercube on Fig. 4(b), we also observe that for $\gamma\leq 1$, the best-performing regimes are $\omega\in[0.5,1]$ as their entropy decreases most rapidly. From $\gamma=2$ onwards, we observe that the entropy of the quantum regime decreases most rapidly, and this rate of decay increases as we move on to $\gamma=4$. For $\gamma=50$ we see that the quantum regime still has the fastest decaying entropy, but this decay is slower than for $\gamma=4$, for example, as performance is less good, as can be seen in Fig. 2. Moreover, when $\gamma=50$ we observe that the purely classical regime, i.e. $\omega=1$, has the slowest decaying entropy after $\omega=0.1$ and $\omega=0.2$. For the cycle graph case, we can see from Fig. 4(c) that when $\gamma=15$ the quantum regime is still less efficient than most hybrid regimes and the classical case. This graph requires higher values of $\gamma$ to accelerate the rate of entropy decay than the complete and hypercube graphs. Finally, for the lollipop where we mark the path vertex, we observe that an increase in $\gamma$ has little impact on the rate of entropy decay. In particular, we see that the quantum regime does not reach its entropy maximum and thus does not start to decay entropy even for high values of the $\gamma$ parameter. The hybrid regime, on the other hand, is affected, albeit only slightly. Looking at Fig. 2, we see that for this graph, when searching for the path vertex, the increase in $\gamma$ only starts to have an effect at $\gamma=15$, reaching a maximum of just 40% transfer efficiency.
@@
(d) Lollipop graph path vertex
Figure 4: Von Neumann entropy $S(\rho)$ as a function of time $t$ for different values of $\gamma$ for the complete graph $K_{64}$, the 6-hypercube $Q_{6}$, the cycle graph $C_{64}$ and the lollipop graph $L_{32,32}$ for the path vertex. The markers indicate the time $t_{S}$ at which the entropy reaches its maximum value before decreasing down to zero.
We also illustrate the reduction of duration $t_{S}$ required to reach the maximum entropy in Fig. 5 for these graphs. For the complete graph, we can clearly see that this duration $t_{S}$ is minimal for low values of $\gamma$ and then increases again as we increase the value of $\gamma$, as an increase in this parameter reduces transfer efficiency. This phenomenon is particularly marked for $\omega\in[0,0.1]$. We point out that the evolution of $t_{S}$ is constant for the classical case for all graphs, i.e. $\omega=1$, as the parameter $\gamma$ lies in the Hamiltonian. We observe much the same behavior in the evolution of time $t_{S}$ for the hypercube, with a sharp decrease followed by an increase as $\gamma$ increases. For both graphs, the best value of $\gamma$ is low, so higher values increase the time needed to reach the entropy maximum before the entropy decreases. For the cycle and the lollipop graphs, we observe a monotonic decay of $t_{S}$ as $\gamma$ increases because for these graphs, an increase in $\gamma$ translates into an improvement in transfer efficiency. In addition, for the lollipop graph, the duration $t_{S}$ for the quantum regime is constant because, as mentioned above, entropy does not reach its maximum with the evolution time we have used. Therefore, given a value of $\gamma$, we observe that the optimal interpolation regime $\omega$ in our framework is the one that leads to the fastest convergence to zero entropy. Furthermore, an increase in the value of $\gamma$ can improve the SQWS performance if it results in a significant reduction in the time needed to reach maximum entropy.
Figure 5: Evolution of the duration $t_{S}$ needed to reach maximum entropy as a function of $\gamma$ for the complete graph $K_{64}$, the 6-hypercube $Q_{6}$, the cycle graph $C_{64}$ and the lollipop graph $L_{32,32}$ (with path vertex) for different values of interpolation $\omega\in[0,1]$.
@@ IV Conclusion
In summary, we have studied a continuous-time search problem in which the walker explores a graph according to a searching dynamics driven to a specific target vertex. We also connected a trapping sink vertex to the target with an irreversible transition. We were interested in the transfer efficiency of a slightly modified search problem, and not in the its computational complexity of the search problem. We extended the result of Caruso et al. [53, 54] to a Stochastic Quantum Walk Search (SQWS) and we numerically showed that a tunable mixing of unitary and non-unitary dynamics can lead to a better transfer efficiency than a non-hybrid evolution for this problem depending on (i) the graph topology, (ii) the target vertex connectivity and (iii) a parametrized Hamiltonian. In particular, the Hamiltonian parameter controls the strength of the quantum walk exploration, its increase reduces or even eliminates the impact of the oracle and therefore the search feature. We have also related the optimal tunable mixing of unitary and non-unitary operations to the system entropy decay rate. Moreover, we have shown that the hybrid regime can beat the purely quantum dynamics only in the presence of a trapping sink. For numerous graphs, mostly sparse, quantum evolution requires to increase the Hamiltonian parameter to achieve the same performance. Therefore, by considering the value of this parameter as a computational resource, the hybrid evolution may require fewer resources than the quantum to perform. More importantly, at a fixed parameters configuration, we can still play with the interaction graph to fit the optimal transfer performance. This can pave the way to a technological leap where noise can be seen as a useful physical resource, in a hardware setting where one can change the connectivity of the architecture for reliable quantum computing. In conclusion, future work could also focus to provide a natively circuit based model for the above results, introducing quantum noise as close as possible to real physical devices’.
@@ Appendix A SQWS with no sink
In this section, we briefly discuss the performances of the SQWS with no use of an extra sink vertex connected to the target vertex. We set $\Gamma=0$ in Eq. (9). This setting corresponds to the usual quantum search problem when $\omega=0$ as the success probability is now related to the presence of the walker in the target vertex $m$ instead of the sink. We use Simulated Annealing [64] to find the best value $\gamma^{*}$ that maximizes the usual spatial search success probability:
$P(\omega,\gamma,t)=\text{Tr}\left({\mathcal{E}[\rho(t)]\ket{m}\bra{m}}\right).$ (13)
Childs and Goldstone search provides a quadratic speedup for the complete graph and the hypercube [42], meaning that the evolved state $\mathcal{E}[\rho(t)]$ reaches a large overlap with the target |$\ket{m}\bra{m}$ | in a timeframe that scales as $t=\mathcal{O}(\sqrt{|V|})$, which is not the case for the cycle graph [62]. Therefore, we run the SQWS on instances of the complete graph, the hypercube and the cycle graph and show the results in Fig. 6. Unsurprisingly, the introduction of non-unitary operations, i.e. increasing the value of $\omega$, drastically reduces search performance. For the complete graph and the hypercube, we still observe a peak reached in a time scaling quadratically with their size when $\omega\leq 0.4$. However, the maximum amplitude reached is much lower than that obtained for the purely unitary evolution, i.e. $\omega=0$. Furthermore, for the complete graph and the hypercube, once $\omega>0.4$, we observe a change in behavior. As there is no longer any oscillatory behavior and we get that $\gamma^{*}=0$, the success probability suddenly converges to 1. However, the time taken to approach 1 no longer scales quadratically with the size of the graph. As for the cycle graph, it is clear that the search does not work even for the quantum case. There is a slight oscillation as long as $\omega\leq 0.3$, then a jump in performance from $\omega=0.4$. However, when $\omega\in[0.4,1]$, although the maximum amplitude reached continues to increase, it follows a very slow growth rate, not exceeding 0.3 when those of the complete graph and the hypercube are very close to 1.
Therefore, we can see that simply removing the sink term completely changes the reaction of the SQWS to the introduction of non-unitary operations. In the presence of a sink, the transfer to the sink can be much more efficient in a noisy regime than in a purely quantum one, based on the graph topology, the target vertex connectivity, and the value of $\gamma$. Hence, the addition of non-unitary operations is not usefull if the graph does not have an extra trapping site.
@@
Figure 6: Probability $P(\omega,\gamma^{*},t)$ over time of finding the target vertex using the Stochastic Quantum Walk Search (SQWS) with no sink vertex connected to the target vertex, i.e. $\Gamma=0$ in Eq. (9), on the complete graph ($K_{64}$), the 6-dimensional hypercube ($Q_{6}$) and the cycle graph ($C_{64}$). The usual quantum walk search is recovered for $\omega=0$, the classical random walk search for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
@@ B.1 Additional graphs
We now run the SQWS exclusively on non-vertex transitive graphs, i.e. graphs with a structure that distinguishes their vertices, of sizes $N\in[63,81]$. We select the path graph ($P_{N}$), the star graph ($S_{N-1}$), the wheel graph ($W_{N}$), the 2D-grid ($G_{\sqrt{N}\times\sqrt{N}}$), the perfect binary tree of depth $d$ ($PBT_{d}$), a maze graph [^4] Maze generation can be easily done using Depth-First Search (DFS) on a grid. Once the maze is created, each cell is considered as a vertex, and two vertices are adjacent if there is no wall between their respective cells. ($M_{N}$) and a random graph ($SW_{N}$) constructed by gluing together three small-world graphs of 22 vertices each with different average connectivity and rewiring probabilities [67]. We show the graph $SW_{66}$ in Fig. 8 and the results of the SQWS on these graphs in Fig. 7. For the maze we mark the exit vertex and for the path, the center and one of the two end vertices, which we call border. The behavior of the maze and path graphs is similar to that of the cycle observed in Fig. 2. The hybrid regime outperforms the quantum one up to a certain value of $\gamma$. Moreover, we note that this critical value, where quantum outperforms hybrid, is highest when searching for a vertex located at the path extremity. We also note that this vertex has the highest eccentricity of all the vertices on which we have run the SQWS (see Table 1). For these two graphs, an increase in $\gamma$ compensates for the increase in $\omega$, as the hybrid regime performs almost as well as the quantum one for high values of $\gamma$. For the star and wheel graphs, unsurprisingly, we obtain very different behavior depending on the marked vertex. When searching for the central vertex, the transfer efficiency is very high from $\gamma=0.05$ except for $\omega\in[0,0.2]$. Indeed, for $\omega>0.2$ we already obtain an efficiency of over 70% for both graphs. The $\omega\in[0,0.2]$ regimes exceed 70% efficiency from $\gamma=1$ and the quantum regime also becomes the best for $\gamma=1$. It can be seen that both graphs react similarly to an increase in $\gamma$ to compensate for the increase in $\omega$. For both graphs, the increase in $\gamma$ ends up slightly reducing performance for $\omega\in[0,0.8]$, and the wheel graph is slightly more affected by this reduction in performance. However, if we do not search for the central vertex, the behavior of the SQWS changes completely. This time, a much higher value of $\gamma$ is required to achieve an efficiency of around 90%. Above a certain value of $\gamma$, approximatively 30 for the star and 15 for the wheel, the quantum regime outperforms the hybrid. We also observe that it is more difficult to stabilize hybrid performance with the increase of $\omega$ for the star than for the wheel graph. Finally, the behavior of the SQWS on the remaining graphs, i.e. the 2D-grid, the perfect binary tree and the random graph composed of three small-world graphs, are quite similar. In each case, the quantum regime becomes more efficient than the hybrid for low values of $\gamma<10$. We note, however, that the grid is the graph for which the hybrid best compensates for the increase in non-unitary dynamics by the increase in $\gamma$, especially when the marked vertex is at one of the extremities. This compensation is least present for the binary tree when searching for the vertex located at depth 5, i.e. the leaf. In this case, the hybrid regime struggles to catch up with the performance of the quantum one once it has passed it. Furthermore, it is for the search of this vertex that the quantum regime requires the highest value of $\gamma=15$ to exceed 80% efficiency among these last three graphs. Moreover, for the root search, we observe one of the same phenomenon as for the complete graph, i.e. that a low-noise regime is less efficient than a noisier one. Finally, the search on the random graph gives rather equivalent results for the three vertices searched. The quantum regime exceeds 80% efficiency for very low values of $\gamma<6$, then increasing this same parameter compensates for the increase in $\omega$. As in the case of the binary tree, we can see that the higher the connectivity of the vertex we are looking for, the more efficient the high-noise regime is compared to the low-noise one, although the difference is very small compared to what we observed for the complete graph.
@@
Figure 7: Stochastic Quantum Walk Search (SQWS) performances with different values of $\gamma\in[0.05,50]$ on a maze graph ($M_{73}$), the path graph ($P_{65}$), the star graph ($S_{63}$), the wheel graph ($W_{64}$), the 2D-grid ($G_{9\times 9}$), the perfect binary tree of depth 5 ($PBT_{5}$) and a random graph ($SW_{66}$) which was constructed by gluing together three small-world graphs of size $N=22$ each with different average connectivity and rewiring probabilities. The time evolution is set to $t=10N$ units of time, with $N$ the size of each graph. The quantum walk dynamics is recovered for $\omega=0$, the classical random walk for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
@@
Figure 8: Random graph $SW_{66}$ constructed by gluing together three small-world graphs of size $N=22$ each with different average connectivity and rewiring probabilities. The target vertices are named HC, IC and LC, respectively for high, intermediate and low connectivity. Their average connectivy are respectively 10, 4 and 3. Their respective rewiring probabilities are 0.1, 0.5 and 0.8.

### v3

@@ III.1.2 Numerical results
Tadpole $T_{M,N}$ MM+N=32N\!=\!32+3232 0.0317 cycle 0.0317 $48\!=\!N+\lfloor M/2\rfloor=\mathcal{O}(N$+$M)$
@@
Maze $M_{N}$ $N=73$ 0.0273 exit 0.0138 34
Figure 3: Instances $T_{8,8}$ and $L_{8,8}$ of the tadpole (upper) and lollipop (lower) graphs, that are respectively a fusion between the cycle graph $C_{8}$ or the complete graph $K_{8}$ with the path graph $P_{8}$. We respectively refer to the locations of the red, green and blue vertices as cycle, shared and path vertices for the tadpole graph, and complete, shared and path for the lollipop graph.
@@ III.2 Relation to entropy
We now relate the performance of an interpolation regime $\omega$ to the evolution of the Von Neumann entropy of the system:
@@
We observe for every run of the SQWS that the entropy first increases up to a maximum, and then decreases until it converges to zero. We also observe that when an increase in the value of $\gamma$ increases the performance of the SQWS, i.e. the transfer to the sink, this translates into a reduction in the time needed to reach the entropy maximum and a faster convergence to zero thereafter. Furthermore, we observe that the interpolation regime $\omega$ whose entropy converges to zero the fastest has the highest transfer efficiency. The presence of the sink vertex $\phi$ introduces dissipation in the system, therefore, as $t\rightarrow\infty$ the walker will end up in the sink with a probability of 1. Thus, the state of the walker converges from $\rho(0)$ to the projector $|\ket{\phi}\bra{\phi}$. Although zero entropy indicates a pure state, it does not guarantee that this state is the state towards which the system converges. Therefore, we also compute the $l_{1}$-norm coherence, which is the sum of the off-diagonal elements of $\rho$:
@@
As the initial state is the uniform superposition over the set of vertices $V$, the initial value is $C_{l_{1}}(\rho(0))=|V|-1$. For all graphs on which we run the SQWS, we observe that this quantity decreases monotonically and it converges to zero. Thus, the convergence of entropy to zero does indicate that the walker state is getting closer and closer to the state $|\ket{\phi}\bra{\phi}$, as $\rho$ becomes increasingly diagonal over time. We illustrate the entropy evolution over time in Fig. 4 for the complete, the hypercube, the cycle and the tadpole (for the cycle vertex) graphs.
Looking at Fig. 4(a), we see that for the complete graph the regime with entropy converging most rapidly to zero for $\gamma=0$ is $\omega=1$. Indeed, the random walk search achieves a transfer efficiency of 89% compared to 87% for $\gamma=0$. We then see that from $\gamma=0.5$ onwards, the quantum regime converges more quickly, as it achieves an efficiency of 94%, then converges even more quickly for $\gamma=1$, where it reaches its maximum efficiency of 97%. We also observe that from $\gamma=2$ onwards, convergence towards zero becomes slower, then when $\gamma$ continues to increase, the chosen evolution time is not sufficient for the entropy to reach its maximum and begin to converge towards zero, which results in very poor performance. It is also clear that the $\omega=0.1$ regime performs less well than the others due to its slow convergence. For the hypercube, we see on Fig. 4(b) that the quantum regime is the most efficient for $\gamma=0$, and is already less efficient than the classical one for $\gamma=1$ as it can be seen on Fig. 2. From $\gamma=2$ onwards, we see that the entropy of the quantum regime no longer reaches its maximum with the chosen evolution time, and the most efficient regimes are those with the highest values of $\omega$. As expected, we observe different behaviours for the cycle graph and the cycle vertex of the tadpole graph  [^6] Note that the same results occur for the maze, the path and the other vertices of the tadpole, but we only display two of them. on Figs. 4(c) and 4(d). This time, we see that the weakly noisy regime is the most effective for low values of $\gamma$, while the other regimes, except for the quantum regime, see their entropy decrease more rapidly with increasing $\gamma$ and eventually decrease more rapidly than $\omega=0.1$. Finally, we observe that entropy converges more slowly than for the complete graph and the hypercube, which explains the lower maximum transfer efficiency levels achieved.
@@
(d) Tadpole graph cycle vertex
Figure 4: Von Neumann entropy $S(\rho)$ as a function of time $t$ for different values of $\gamma$ for the complete graph $K_{64}$, the 6-hypercube $Q_{6}$, the cycle graph $C_{64}$ and the tadpole graph $T_{32,32}$ for the cycle vertex. The markers indicate the time $t_{S}$ at which the entropy reaches its maximum value before decreasing down to zero.
We also illustrate the reduction of duration $t_{S}$ required to reach the maximum entropy in Fig. 5 for these graphs. We can see that for the complete graph and the hypercube, variations in $\gamma$ mainly affect the duration $t_{S}$ for $\omega=0$ and $\omega=0.1$. The most sensitive regime is the quantum regime, and we observe that the duration $t_{S}$ reaches its minimum for $\gamma=1$ and $\gamma=0.43$, respectively, which correspond to the values of $\gamma$ that give the highest transfer efficiency for these graphs. We also see that once these values are exceeded, this duration increases dramatically, which delays the moment when entropy will decrease to begin converging towards zero and thus results in a significant loss of performance. For the cycle and tadpole graphs, variations in $\gamma$ have no effect on the quantum regime and seem to affect only the low-noise regime for $\omega=0.1$ and $\omega=0.2$.
Figure 5: Evolution of the duration $t_{S}$ needed to reach maximum entropy as a function of $\gamma$ for the complete graph $K_{64}$, the 6-hypercube $Q_{6}$, the cycle graph $C_{64}$ and the tadpole graph $T_{32,32}$ (with cycle vertex) for different values of interpolation $\omega\in[0,1]$.
@@ IV Conclusion
In summary, we have studied a continuous-time search problem in which the walker explores a graph according to a searching dynamics driven to a specific target vertex. We also connected a trapping sink vertex to the target with an irreversible transition. We were interested in the transfer efficiency of a slightly modified search problem, and not in the its computational complexity of the search problem. We extended the result of Caruso et al. [55, 56] to a Stochastic Quantum Walk Search (SQWS) and we numerically showed that a tunable mixing of unitary and non-unitary dynamics can lead to a better transfer efficiency than a non-hybrid evolution for this problem depending on (i) the graph topology, (ii) the target vertex connectivity and (iii) a parametrized Hamiltonian. In particular, the Hamiltonian parameter controls the strength of the oracle that marks the target vertex. We have also related performance of an interpolation regime to the system entropy decay rate. Moreover, we have shown that the hybrid regime can beat the purely quantum dynamics only in the presence of a trapping sink. For numerous graphs, mostly dense, quantum evolution is very sensitive to an increase in the strength of the oracle which is not the case for the hybrid regime. We have also shown that low-noise regimes give the worst performance for most of the graphs, expect for sparse graphs, with low degree centrality and high eccentricity for the target vertex, where these regimes are the most efficient when the strength of the oracle remains fairly weak. Therefore, by considering the value of this parameter as a computational resource, the hybrid evolution may require fewer resources than the quantum to perform on specific topologies. More importantly, at a fixed parameters configuration, we can still play with the interaction graph to fit the optimal transfer performance.
This can pave the way to a technological leap where noise can be seen as a useful physical resource, in a hardware setting where one can change the connectivity of the architecture for reliable quantum computing. In conclusion, future work could also focus to provide a natively circuit based model for the above results, introducing quantum noise as close as possible to real physical devices’.
@@ Appendix A SQWS with no sink
In this section, we briefly discuss the performances of the SQWS with no use of an extra sink vertex connected to the target vertex. We set $\Gamma=0$ in Eq. (9). The success probability is now related to the presence of the walker in the target vertex $m$ instead of the sink:
[…]
Childs and Goldstone search provides a quadratic speedup for the complete graph and the hypercube [44], meaning that the evolved state $\mathcal{E}[\rho(t)]$ reaches a large overlap with the target $|\ket{m}\bra{m}$ in a timeframe that scales as $t=\mathcal{O}(\sqrt{|V|})$, which is not the case for the cycle graph [64]. Therefore, we run the SQWS on instances of the complete graph, the hypercube and the cycle graph and respectively show the results in Figs. 6, 7 and 8. For the complete graph, we find an oscillatory behaviour for the probability of success when $\omega=0$ because the dynamics are completely unitary. We see that the highest performance is obtained for $\gamma\in[0.8,1.5].$ Then, as soon as $\omega$ increases, we lose the oscillatory behaviour and the quantum regime no longer performs well. Furthermore, for the hybrid regime, we see that we obtain a probability of success of 0.8 for much longer evolution times and higher values of $\gamma$ than were necessary to obtain the same probability with the quantum regime. When $\omega=0.8$, we obtain a probability of success of approximately 0.7 for values of $\gamma<1$. Finally, we obtain almost similar results for $\omega=0.9$ and $\omega=1$, although the time required to reach a 0.8 probability of success is much higher than that required for the quantum regime. Similar results are obtained for the hypercube, except that the optimal values of $\gamma$ are lower than for the complete graph for the quantum regime. Otherwise, for the hybrid regime, we observe more or less the same results, and we must wait for the classical regime $\omega=1$ to achieve a probability of success of approximately 0.8 for values of $\gamma<1$, although the evolution time required is much longer than for the quantum regime. For the graph cycle, we obtain different results for the quantum regime. The highest probability of success is 0.4 and is achieved for a considerably higher time value than for the two previous graphs. For the hybrid and classical regimes, we obtain results similar to the previous graphs, although an even longer evolution time is required to reach the maximum probability of success of approximately 0.4.
These results indicate that the presence of the sink is necessary to observe phenomena where the hybrid case outperforms the quantum regime, because without its presence, the addition of non-unitary dynamics only degrades performance.
@@
Figure 6: Probability $P(\omega,\gamma,t)$ over time of finding the target vertex using the Stochastic Quantum Walk Search (SQWS) with no sink vertex connected to the target vertex, i.e. $\Gamma=0$ in Eq. (9), on the complete graph ($K_{64}$). The quantum walk search is recovered for $\omega=0$, the classical random walk search for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
Figure 7: Probability $P(\omega,\gamma,t)$ over time of finding the target vertex using the Stochastic Quantum Walk Search (SQWS) with no sink vertex connected to the target vertex, i.e. $\Gamma=0$ in Eq. (9), on the hypercube graph ($Q_{6}$). The quantum walk search is recovered for $\omega=0$, the classical random walk search for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
Figure 8: Probability $P(\omega,\gamma,t)$ over time of finding the target vertex using the Stochastic Quantum Walk Search (SQWS) with no sink vertex connected to the target vertex, i.e. $\Gamma=0$ in Eq. (9), on the cycle graph ($C_{64}$). The quantum walk search is recovered for $\omega=0$, the classical random walk search for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
@@ B.1 Additional graphs
We now run the SQWS exclusively on non-vertex transitive graphs, i.e. graphs with a structure that distinguishes their vertices, of sizes $N\in[63,81]$. We select the lollipop graph ($L_{M,N}$), the star graph ($S_{N-1}$), the wheel graph ($W_{N}$), the 2D-grid ($G_{\sqrt{N}\times\sqrt{N}}$), the perfect binary tree of depth $d$ ($PBT_{d}$) and a random graph ($SW_{N}$) constructed by gluing together three small-world graphs of 22 vertices each with different average connectivity and rewiring probabilities [67]. We show the graph $SW_{66}$ in Fig. 10 and the results of the SQWS on these graphs in Fig. 9. The lollipop consists of two graphs for which SQWS behaves differently, namely the complete graph and the path graph. Similar results are obtained for the complete and shared vertices with a transfer efficiency of 49% for all values of $\omega\neq 0.1$ and when $\gamma<2$. There is a performance drop of about 10% for the $\omega=0.1$ regime compared to the others. Finally, when $\gamma>2$, the quantum regime loses efficiency drastically, reaching 0.02% for $\gamma=30$. When the target vertex is path, however, the results are completely different: SQWS performance is extremely poor for all $\omega$ regimes and does not exceed 0.05%. We also observe that, although low, the efficiency is insensitive to variations in $\gamma$. However, it increases slightly with the increase in $\omega$, from 0.01% for $\omega=0$ to 0.05% for $\omega=1$. For star and wheel graphs, the results are also similar to each other depending on the target vertex. When the target is the central vertex of the graph center, we obtain respective efficiencies greater than 87% and 80% for all regimes $\omega$ when $\gamma\leq 2$. For both graphs, the quantum regime remains slightly more efficient than the others, with 97% for $\gamma=1$. Furthermore, once $\gamma>2$, we also observe a decrease in performance, reaching 0.02% for $\gamma=30$ for both graphs. We obtain very different results when the target vertex is border, which is any vertex other than the central vertex. For both graphs, variations in $\gamma$ do not impact performance, and efficiency increases with increasing $\omega$. For the star graph, efficiency does not exceed 0.08%, and 0.16% for the wheel graph. Although efficiency is very low in both cases, the results for the wheel graph are twice as good as those for the star graph. For the grid, we mark the central vertex and a vertex located at one end. We obtain twice as good results for centre as for border. Furthermore, for centre, the highest performance is obtained for $\omega=0$ and $\omega=1$, whereas for border, it is for $\omega=1$. For the center vertex, we observe that for the noisy regime, an increase in $\omega$ improves performance, reaching a maximum of around 70% for $\omega=1$. As for the center vertex, the quantum regime is never the most efficient; moreover, we observe that for low values of $\gamma\leq 0.3$, the low-noise regime $\omega=0.1$ performs better than the regimes $\omega\in[0.2,0.8]$, which we only observed for graphs with low density and centrality degree and high eccentricity. Of all the vertices on which we ran SQWS, the border vertex of the grid most closely approximates these characteristics apart from those of the cycle, path, tadpole and maze, which explains the weak appearance of this phenomenon. For the perfect binary tree with depth 5, we mark three vertices located at different depths in the tree: root (depth 0), a vertex depth3 (located at depth 3), and a leaf vertex at depth 5. Overall, we observe that the greater the depth of the target vertex, the poorer the performance of SQWS. Finally, for the random graph shown in Fig. 10 composed of different graphs with different densities, we observe the same phenomenon, where performance decreases with the loss of density and connectivity of the target vertex.
@@
Figure 9: Stochastic Quantum Walk Search (SQWS) performances with different values of $\gamma\in[0,30]$ on the lollipop graph ($L_{32,32}$), the star graph ($S_{63}$), the wheel graph ($W_{64}$), the 2D-grid ($G_{9\times 9}$), the perfect binary tree of depth 5 ($PBT_{5}$) and a random graph ($SW_{66}$) which was constructed by gluing together three small-world graphs of size $N=22$ each with different average connectivity and rewiring probabilities. The time evolution is set to $t=10N$ units of time, with $N$ the size of each graph. The quantum walk dynamics is recovered for $\omega=0$, the classical random walk for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
@@
Figure 10: Random graph $SW_{66}$ constructed by gluing together three small-world graphs of size $N=22$ each with different average connectivity and rewiring probabilities. The target vertices are named HC, IC and LC, respectively for high, intermediate and low connectivity. Their average connectivity are respectively 10, 4 and 3. Their respective rewiring probabilities are 0.1, 0.5 and 0.8.